OpenStudy (anonymous):

Show that fx(0, 0) and fy(0, 0) both exist, but that f is not differentiable at (0, 0) where f define as f(x, y) = { (-3xy)/(x^2 + y^2) if (x, y) ≠ 0 0 if (x, y) = (0, 0)

OpenStudy (zehanz):

By fx and fy, do you mean the partial derivatives with respect to x and y?

OpenStudy (anonymous):

yes

OpenStudy (zehanz):

The partial derivatives are:\[\frac{ \partial f}{ \partial x }(x, y)=-\frac{ 3y(x^2+y^2)-3xy2x }{ (x^2+y^2)^2 }=...=\frac{ 3y(x^2-y^2) }{ (x^2+y^2)^2 }\]\[\frac{ \partial f}{ \partial y }(x, y)=-\frac{ 3x(x^2+y^2)-3xy2y }{ (x^2+y^2)^2 }=...=\frac{ -3x(x^2-y^2) }{ (x^2+y^2)^2 }\]Now I'm going to do some thinking... ;)

OpenStudy (anonymous):

@ZeHanz excuse me!! I think you got confuse with my question -.-

OpenStudy (anonymous):

\[f _{x}(0,0)=Lt _{h \rightarrow0}f((0+h,0)-f(0,0))/h\]

OpenStudy (anonymous):

so........after substitution u get this \[f _{x}(0,0)=0\]

OpenStudy (zehanz):

The partial derivatives I calculated are formulas only valid for (x, y) <> (0, 0). To get the pd's in (0, 0), you only have to calculate:\[\frac{ \partial f }{ \partial x } (0,0)=\lim_{h \rightarrow 0}\frac{ f(0+h, 0)-f(0,0) }{ h }\]and\[\frac{ \partial f }{ \partial y } (0,0)=\lim_{h \rightarrow 0}\frac{ f(0, 0+h)-f(0,0) }{ h }\]The only thing is, with this function f, I don't think that the partial derivatives actually exist in (0,0).

OpenStudy (anonymous):

they do exist are equal to zero.........@ZeHanz

OpenStudy (anonymous):

u know that f(x,y) is differentiable at the origin(0,0) if \[f(h,k)-f(0,0)-hf _{x}(0,0)-kf _{y}(0,0)=h \epsilon _{1}+k \epsilon _{2}\]

OpenStudy (anonymous):

where \[\epsilon _{1}\] and \[\epsilon _{2}\] tend to zero as h,k tend to zero independently

OpenStudy (zehanz):

My fault, the partial derivatives are 0 in (0, 0), I see that now (could use l'H to prove). If f is discontinuous in (0, 0), then it is not differentiable...

OpenStudy (anonymous):

so i have to find f is not continues in (0, 0) then it's both exist right?

OpenStudy (anonymous):

@ZeHanz @nitz ........

OpenStudy (zehanz):

No, we already know that the partial derivatives exist in (0,0). The both have value 0 there. We now have to prove that f is not continuous in (0,0). This means that when (x,y) is chosen near to (0,0), f(x,y) is not near to f(0,0)=0. Because f must at least be continuous in (0,0) to be differentiable there, failing this condition means f is not differentiable in (0,0).

OpenStudy (anonymous):

Do you mind telling me more about the solution for this.

OpenStudy (zehanz):

Because continuity is a necessary condition for differentiability,I hoped to prove f is not continuous in (0,0). Now I'm not sure f isn't continuous in (0,0). It could well be, because it seems to me the limit of f as (x,y) approaches (0,0) is 0, exactly f(0,0). If it is continuous, but not differentiable, the graph has some kind of sharp edge in (0,0). (like in the one-dimensional case the graph of f(x)=|x| has in x=0) The whole thing is rather complex. Here is a PowerPoint about it, I've found on the internet: www2.kenyon.edu/Depts/.../Differentiability.ppt

OpenStudy (zehanz):

Even better: http://higheredbcs.wiley.com/legacy/college/hugheshallett/0471484822/theory/hh_focusontheory_sectioni.pdf See Example 4 in this document. It is about the function f(x,y)=xy/(x²+y²), so (apart from the non-important factor -3) it's just what you are looking for! Your question is answered there.

OpenStudy (zehanz):

It turns out that f is indeed NOT continuous in (0,0), so it is also NOT differentiable, although the partial derivatives exist there...

OpenStudy (zehanz):

Why is it not continuous in (0,0)? To be continuous, this should be true: \[\lim_{(x,y) \rightarrow (0,0)}f(x,y)=f(0,0)=0\]It means, that along whatever path you walk to (0,0), you should always arrive at 0, which is the value of f in (0, 0). But: if you walk along the path (x,x), i.e. along the line y=x, then the limit becomes:\[\lim_{x \rightarrow 0}\frac{ -3x^2 }{ 2x^2 }=-\frac{ 3 }{ 2 }\neq0\]So it doesn't lead to f(0,0) = 0, but to -3/2. Conclusion: f not continuous in (0,0), so also not differentiable.

OpenStudy (anonymous):

@ZeHanz THank you

OpenStudy (zehanz):

YW!